# Detecting Harmful Memes with Decoupled Understanding and Guided CoT Reasoning

### v1

@@ 5 Conclusion
In this paper, we focus on the task of harmful meme detection and address the limitations of existing methods in resource efficiency, flexibility, and explainability. We introduce U-CoT+, a framework that converts the default multimodal setting into a unimodal, text-only task using a high-fidelity meme-to-text pipeline, and further boosts the zero-shot performance of LLMs through a novel guided CoT prompting scheme. By incorporating high-quality, targeted, and interpretable human-crafted guidelines, U-CoT+ enables small-scale LLMs to achieve strong performance comparable to fully supervised baselines and, to match or outperform existing low-resource baselines and advanced GPT-4-based models across seven datasets. These results demonstrate the effectiveness of our framework for low-resource, explainable harmful meme detection.
Limitations
We consider the following limitations for future work. First, we focus solely on the zero-shot harmful meme classification performance of individual models i.e., we report results based only on each LLM’s independent predictions. We may later explore whether ensemble methods, such as majority voting across the decisions given by multiple small-scale LLMs and LMMs, can further improve zero-shot performance. Second, we currently only rely on LLMs’ internalized knowledge for zero-shot classification, it is yet to explored whether small-scale LLMs under our proposed guided CoT prompting can further benefit from Retrieval Augmented Generation Anaissi et al. (2025) that equips LLMs with additional knowledge.
Ethics Statement
This research exclusively utilizes publicly available datasets and does not involve any personally identifiable information. We have taken steps to minimize potential biases by implementing explicit, human-crafted guidelines to guide unimodal LLMs’ behaviors in detecting harmful content in memes. Our approach prioritizes transparency and interpretability, enabling human oversight of model predictions and reasoning steps. All code and data used in our experiments are released to support reproducibility and community validation. We acknowledge that some readers may find certain content disturbing, and we advise appropriate caution when reviewing examples included for demonstration purposes.
Appendix A Datasets
We experiment with the following datasets:
Datasets Test Train
#Harm #Benign
FHM Kiela et al. (2020) 490 510 8500
HarMeme Pramanick et al. (2021a) 124 230 3013
Harm-P Pramanick et al. (2021b) 171 184 2939
MultiOFF Suryawanshi et al. (2020) 58 91 445
MAMI Fersini et al. (2022) 500 500 9000
PrideMM Shah et al. (2024) 247 260 4328
GB(Hateful) Lin et al. (2024b) 750 1250 -
GB(Harmful) 420 589 -
GB(Offensive) 303 440 -
GB(Misogynistic) 500 500 -
Table 5: Dataset Statistics. GB: The Goat-Bench dataset Lin et al. (2024b). The sizes of the training splits are provided as a reference for the training data scale used in SFT-based models.
  • FHM Kiela et al. (2020) The Facebook Hateful Memes (FHM) dataset was released in 2021 as a challenge set for multimodal classification, focusing on detecting hate speech in multimodal memes. The dataset was synthetically constructed to examine model performance in distinguishing real hateful memes from harmless ones that include their benign confounders, which are constructed with flipped labels through minimal changes to one of the modalities (either images or overlaid text). The hatefulness of a meme can be determined by either both or only one of the modalities Wu et al. (2020, 2022, 2023b, 2023a, 2024c, 2024d, 2024b, 2024a). Hateful speech under the context of this dataset is defined as: A direct or indirect attack on people based on characteristics, including ethnicity, race, nationality, immigration status, religion, caste, sex, gender identity, sexual orientation, and disability or disease. We define attack as violent or dehumanizing (comparing people to non-human things, e.g. animals) speech, statements of inferiority, and calls for exclusion or segregation. Mocking hate crime is also considered hate speech.
  • HarMeme Pramanick et al. (2021a) contains 3,544 online memes related to COVID-19 pandemic. They define “harmful” in a broader way different from “hateful” or “offensive” as the potential to cause “mental abuse, defamation, psycho-physiological injury, proprietary damage, emotional disturbance, and compensated public image”. The targets of these harmful memes can be any individual, organization, community, or the general society. Given the inherently ambiguous nature of such definition, which could be different from general criteria internalized in most LLMs, we find that annotated labels in this dataset are highly biased to favor specific types of memes as harmful, particularly those that may harm the public image of Donald Trump.
  • HarmP Pramanick et al. (2021b) is a follow-up work of HarMeme but has a shift in attention to a broader collection of U.S. political memes, which are labelled into three categories i.e., “very harmful”, “somewhat harmful” and “not harmful”, following the same annotation procedure in HarMeme. Following prior works Mei et al. (2025), we incorporate “very” and “somewhat” harmful memes into a unified “harmful” class. Yet, the ambiguity in “harmfulness” definition may have resulted in unclear criteria and inconsistent annotation in the data. In addition, unlike synthetic memes that usually follow a standard creation format, online political memes can be highly free-form or noisy, e.g., digitally altered, distorted, masked, blurred, filled with hardly identifiable text etc. Such low-quality images pose great challenges on existing small LMMs.
  • MultiOFF Suryawanshi et al. (2020) is a small collection of political memes similar to those in Harm-P but are mainly related to 2016 U.S. presidential election in terms of context. This datasets highlights a specific taxonomy of multiple types of offensive memes, including “racial abuse”, “attacking minorities”, “personal attacks” and “homophobic abuse”.
  • MAMI Fersini et al. (2022) specifically focuses on misogynistic memes that cause harm to women. Misogynistic memes are generally more explicit in language and imagery compared to other types of harmful memes.
  • PrideMM Shah et al. (2024) contains 5,063 online memes in the context of LGBTQ+ pride movements. Each meme is annotated with labels indicating its stance toward the goal of LGBTQ+ movements, whether it is considered “hateful” or “humorous”. Apart from the binary labels for hatefulness, the dataset defines a taxonomy of hate targets: “Undirected”, “Individual”, “Community” and “Organization”. Whether a meme can be regarded as “hateful” varies with respect to the perspective of different targets. Therefore, this dataset also includes a target classification task.
  • GoatBench Lin et al. (2024b) is a comprehensive meme benchmark comprising over 6K varied memes encapsulating themes such as implicit hate speech, sexism, and cyberbullying, etc., curated from the above datasets. In this work, we only evaluate on their harmfulness-related sub-tasks, without considering the task of Sarcasm.
Appendix B Baselines
We summarize the key techniques and backbone models of the baselines as follows:
  • LMM-RGCL Mei et al. (2025) learns hate-aware vision and language representations through a contrastive learning objective applied to a pre-trained CLIP encoder, achieving state-of-the-art performance on the MultiOFF dataset.
  • UMR Yang et al. (2024) integrates uncertainty-guided modal rebalance with Gaussian-based stochastic embeddings and improved cosine loss on frozen CLIP/ALBEF/BLIP/BLIP-2 backbones to handle modality imbalance in multimodal hate detection.
  • Pro-Cap Cao et al. (2023a) leverages frozen BLIP-2 for zero-shot VQA to generate probing-based captions, which are then combined with meme texts for training BERT and PromptHate models on hateful meme detection.
  • CapAlign Ji et al. (2024) uses LLaMA-2-7B as the base model and fine-tunes it with LoRA on instruction-following and synthetic alignment conversations, enabling efficient and aligned caption generation for multimodal tasks.
  • MemeCLIP Shah et al. (2024) builds on frozen CLIP ViT-L/14 encoders and introduces lightweight linear projections, feature adapters, and a cosine classifier with semantic-aware initialization for multi-aspect meme classification.
  • M3H-p-CoT Kumari et al. (2024) proposes a multi-hop reasoning framework combining EOR features and meme text to prompt a CLIP-based classifier through hierarchical CoT modules trained with cross-attention and contrastive learning.
  • ExplainHM Lin et al. (2024a) fine-tunes LLaVA Liu et al. (2023b) to generate multimodal rationales under debate-style prompts and then trains a T5-based model with multimodal fusion to assess meme harmfulness.
  • IntMeme Hee and Lee (2025) uses frozen InstructBLIP and mPLUG-Owl to generate meme interpretations in a zero-shot setting, then trains RoBERTa and FLAVA encoders on meme content and interpretations for hateful meme classification.
  • LoReHM Huang et al. (2024) adopts an agent-based LMM framework (experimenting on Llava-34B and GPT-4o) that leverages few-shot in-context learning and self-improvement capabilities for low-resource hateful meme detection.
  • Mod-HATE Cao et al. (2024) trains a suite of LoRA modules and utilizes few-shot demonstration examples to train a module composer, which assigns weights to the LoRA modules for effective low-resource hateful meme detection.
  • GPT-4o Hurst et al. (2024) is a cutting-edge multimodal model.
LLM LMM Scheme FHM HarMeme Harm-P MultiOFF MAMI PrideMM
Acc F1 Acc F1 Acc F1 Acc F1 Acc F1 Acc F1
GPT-4o-mini 67.60 65.51 70.90 69.46 65.35 65.35 65.77 64.86 77.40 76.59 72.39 72.28
Llama3.1-8Bf Qwen2VL-7B U-CoT 63.50 61.44 62.43 50.64 $\uparrow$62.25 $\uparrow$58.78 63.76 57.55 $\uparrow$69.80 $\uparrow$69.09 61.34 57.90
LLaVa1.6-7B $\uparrow$63.00 $\uparrow$60.36 65.25 46.66 $\uparrow$59.72 $\uparrow$56.40 $\uparrow$63.09 51.84 $\uparrow$68.70 $\uparrow$68.10 $\uparrow$60.55 58.49
LLaVa+Qwen $\uparrow$62.80 $\uparrow$60.15
Qwen2VL-7B U-CoT+ $\uparrow$68.00 $\uparrow$66.98 $\uparrow$71.19 $\uparrow$61.44 61.13 56.21 $\uparrow$68.46 $\uparrow$62.35 $\uparrow$71.90 $\uparrow$71.86 $\uparrow$63.51 $\uparrow$60.86
LLaVa1.6-7B $\uparrow$67.00 $\uparrow$65.51 $\uparrow$72.03 $\uparrow$60.88 57.75 51.09 $\uparrow$63.76 $\uparrow$53.86 $\uparrow$73.50 $\uparrow$73.50 $\uparrow$64.50 $\uparrow$63.26
LLaVa+Qwen $\uparrow$65.70 $\uparrow$64.10
Qwen2.5-7Bf Qwen2VL-7B U-CoT $\uparrow$66.70 $\uparrow$66.45 48.59 47.07 $\uparrow$58.59 $\uparrow$57.46 57.05 56.81 $\uparrow$71.80 $\uparrow$71.80 63.51 62.38
LLaVa1.6-7B $\uparrow$62.70 $\uparrow$61.94 59.89 59.82 $\uparrow$58.03 $\uparrow$57.80 $\uparrow$63.76 $\uparrow$63.72 $\uparrow$70.60 $\uparrow$70.59 $\uparrow$63.31 $\uparrow$61.98
LLaVa+Qwen $\uparrow$63.80 $\uparrow$63.23
Qwen2VL-7B U-CoT+ $\uparrow$67.50 $\uparrow$67.50 $\uparrow$70.34 $\uparrow$69.97 $\uparrow$62.54 $\uparrow$60.20 $\uparrow$62.42 $\uparrow$62.13 $\uparrow$73.60 $\uparrow$73.39 $\uparrow$67.26 $\uparrow$66.90
LLaVa1.6-7B $\uparrow$65.70 $\uparrow$65.53 $\uparrow$77.12 $\uparrow$75.26 $\uparrow$59.15 $\uparrow$54.99 $\uparrow$67.11 $\uparrow$66.63 $\uparrow$74.40 $\uparrow$74.40 $\uparrow$64.30 $\uparrow$64.00
LLaVa+Qwen $\uparrow$66.60 $\uparrow$66.55
Table 6: Complete results comparing different prompting schemes for zero-shot harmful meme detection using Llama3.1-8B and Qwen2.5-7B. Performance improvements from multimodal (M-CoT) to unimodal text-only (U-CoT) settings are highlighted by $\uparrow$, while improvements from plain to guided CoT (U-CoT+) are highlighted by $\uparrow$. The best results achieved by each LLM are in bold.
LLM LMM Scheme Hatefulness Harmfulness Misogyny Offensiveness
Acc F1 Acc F1 Acc F1 Acc F1
GPT-4V M-CoT 71.90 71.37 66.90 64.97 81.60 81.57 61.37 60.99
Qwen2.5-14Bf Qwen2VL-7B U-CoT 70.80 69.17 60.16 59.43 77.20 77.20 58.95 58.39
LLaVa1.6-7B 70.15 67.44 62.64 60.51 77.40 77.38 60.97 59.76
LLaVa+Qwen 69.70 67.08
Qwen2VL-7B U-CoT+ 71.25 70.75 71.16 70.69 79.30 79.28 59.49 58.91
LLaVa1.6-7B 71.95 71.05 71.56 70.53 80.10 80.09 61.78 60.90
LLaVa+Qwen 70.80 69.95
Mistral-12Bf Qwen2VL-7B U-CoT 70.2 68.21 61.45 58.93 74.9 74.9 63.53 60.62
LLaVa1.6-7B 68.15 64.45 60.46 56.84 75.5 75.48 60.57 56.22
LLaVa+Qwen 68.2 64.67
Qwen2VL-7B U-CoT+ 71.45 70.43 70.76 68.32 73.8 73.73 60.70 58.86
LLaVa1.6-7B 70.05 68.15 70.66 67.63 75.1 74.83 61.1 58.07
LLaVa+Qwen 71.75 70.15
Table 7: Complete harmful meme detection results on Goat-Bench Lin et al. (2024b).
Appendix C Implementation Details
We run the zero-shot inference experiments on two NVIDIA A6000 48GB GPUs and six NVIDIA A40 48GB GPUs. All LMM-based multimodal inference are conducted with a batch size of 1. Qwen2.5-based LLMs and Mistral are run with a batch size of 16. LLama3.1 is run with a batch size of 32. We obtain GPT-4o-mini inference results using the OpenAI API with the model version “gpt-4o-mini-2024-07-18”. For reproducibility, we adhere to greedy decoding without sampling (with “do_sample”’ set to False and “temperature” as 0) to ensure fully deterministic generation. The “max_new_tokens” is set as 256/356 for LMMs and 1024/1536 for LLMs depending on datasets.
C.1 Visual Information Extraction
We experiment with two widely used open-source LMMs with 7 billion parameters i.e., Llava1.6-Mistral-7B Liu et al. (2023a) and Qwen2-VL-7B Wang et al. (2024). The two models have been extensively pre-trained to gain strong capabilities in image understanding, embedded text recognition and multimodal inference, and are generally accessible under most computational resource conditions.
We identify various types of visual clues that are critical for downstream harmful meme detection and design the questions accordingly to prompt LMMs to extract the information through visual question answering (VQA). In particular, for the synthetic FHM dataset, since the visual style of its images is more like that of regular photographs other than that of authentic online memes, the captions of some memes do not directly relate to the visual contents. Therefore, we explicitly add the instruction “Ignore any overlaid text or caption.”, denoted by [Ignore], after the questions to force LMMs to focus solely on the visual content without being distracted by the overlaid captions. For most binary questions, we adopt this template for a unified output format: “Start your response with “Yes,” or “No,” before giving the explanation.”, denoted by [Fm]. The overlaid text or caption in a meme is denoted by [OCR]. Question templates:
Human
Is there any human subject in the given image? ([Ignore]) [Fm]
#Human
Does the image include more than one human subject? ([Ignore]) [Fm]
Gender
What is/are the gender(s) of the human subject(s) in the image?
Race
What is/are the perceived race(s) of the human subject(s) in the image?
Appearances
What are the distinctive physical appearance characteristics of the human subject(s) in the image?
Disability
Does any of the human subjects in the image have any disability?
Celebrity
(i) Who is/are the human subject(s) in the image? (i) Is any celebrity or historical figure portrayed in the image? If yes, who are they? If no, just output “No.”.[Fm] (i) FHM-specific: a. Does the image portray Adolf Hitler? If yes, output “Yes, Adolf Hitler is portrayed in the image.” Otherwise, if you are not sure about the identity of human subject, just output “No, I can’t tell.” b. Does the image portray Anne Frank, the Jewish girl who hid from the Nazis during World War II? If yes, output “Yes, Anne Frank is portrayed in the image.” Otherwise, if you are not sure about the identity of human subject, just output “No, I can’t tell.”
@@ C.1 Visual Information Extraction
Adult Content
Does the image’s visual content contain adult content?[Ignore][Fm]
Female
Is/Are there any human subject(s) in the image female?[Ignore][Fm]
Sexual
(i) Does the image’s visual content highlight the sexiness of the female’s figure in a way that is sexually provocative?[Ignore][Fm] (i) Does this image highlight sexual body parts of the female subject(s), such as the breast, the hip/buttock, or the genital?[Ignore][Fm] (i) Does/Do the female subject(s) in the image appear to be overweight?[Ignore][Fm] (i) Does/Do the female subject(s) in the image appear to be of large body size (considered as fat)?[Ignore][Fm]

### v2

@@ 7. Conclusion
In this paper, we introduce U-CoT+, a novel framework for low-resource harmful meme detection. With a high-fidelity meme2text pipeline, it decouples meme content recognition from harmfulness analysis and converts the default multimodal task into a unimodal, text-based setting, thus enabling detection based on LLMs’ inherent capabilities. By incorporating targeted, interpretable and easily obtained human-crafted guidelines, U-CoT+ further enhances the zero-shot performance of lightweight LLMs with guided CoT prompting, achieving performance highly competitive with, or even surpassing that of resource-intensive baselines across datasets from diverse sociocultural contexts. Experiment results demonstrate U-CoT+’s effectiveness in offering strong resource efficiency, flexibility, scalability and explainability for harmful meme detection.
8. Acknowledgments
This work is supported by AI-X, Interdisciplinary Graduate Programme, and the College of Computing and Data Science of Nanyang Technological University, Singapore.
Appendix A ETHICS STATEMENT
This work aims to fully harness the inherent capabilities of state-of-the-art LLMs for harmful meme detection under low-resource constraints, contributing to the safety and integrity of online communication, particularly on social media platforms. We position our method as an early-stage support tool for real-world content moderation, designed to assist human moderators by providing LLM-generated verdicts as auxiliary references for human verification. In doing so, our method helps alleviate moderators’ emotional and cognitive burden. We see our work as part of the broader effort to foster a safe, respectful and responsible digital environment.
We exclusively use publicly available datasets in compliance with their usage agreements and release our codebase along with detailed implementation specifics to support reproducibility. Our work aligns with the vision of sustainable and accessible AI by leveraging lightweight open-source models that are deployable under common computational conditions.
While our guidelines are not inherently designed to induce bias, their effectiveness may be constrained by pre-existing biases embedded in the underlying LLMs. To mitigate the risk of unfair moderation arising from such LLM-induced biases, it is crucial that human moderators interpret LLM verdicts with caution. Our method is intended solely for the detection and prevention of harmful memes; any use that promotes, condones, or encourages hate speech or harmful content is strictly prohibited and strongly condemned.
Appendix B IMPLEMENTATION DETAILS
We conduct experiments on two NVIDIA A6000 48GB GPUs. All LMM-based inference are conducted with a batch size of 1. Qwen2.5 and Mistral are run with a batch size of 16. LLama3.1 is run with a batch size of 32. We obtain GPT-4o-mini inference results using the OpenAI API with the model version “gpt-4o-mini-2024-07-18”. For reproducibility, we adhere to greedy decoding without sampling (with “do_sample”’ set to False and “temperature” to 0) to ensure fully deterministic generation. The “max_new_tokens” is set as 256 or 356 for LMMs and 1024 or 1536 for LLMs depending on datasets. Table 5 shows detailed dataset statistics.
All prompt templates and workflows for visual information extraction, information integration, fine-grained target identification (e.g., protected group detection and hateful forms generation for FHM, entity identification and target classification for PrideMM) and guided CoT prompting are released in our codebase. Figure 5 visualizes how our proposed meme2text pipeline configures visual questions for different sociocultural contexts.
For ablation study, we query GPT-4o-mini using fine-grained prompts to obtain detail-preserving meme descriptions. For memes that GPT-4o-mini refuses to process (e.g., 14 in FHM and 332 in MAMI), we fallback to our original descriptions based on 7B LMMs. Table 6 illustrates our ensemble strategies that integrate predictions based on different LMMs as well as from U-CoT and M-CoT. Specifically, by default, if there is at least one positive U-CoT+ prediction, we classify the meme as harmful. For datasets where we have high confidence in the alignment between our guidelines and the true positives, memes with no positive U-CoT+ predictions are classified as harmless, regardless of the predictions made by M-CoT or U-CoT. For guidelines with a medium confidence level, memes are also considered harmless if there is only one positive U-CoT+ prediction while all M-CoT and U-CoT predictions are negative. When the guidelines are prone to over-censorship (i.e., with a low confidence level), memes with all negative M-CoT and U-CoT predictions are treated as harmless, regardless of the guideline-based results.
Our error analysis is conducted on a set of error cases with a balanced distribution of false negatives and false positives, randomly sampled from each of the three underperforming datasets (FHM, Harm-P, and PrideMM).
Appendix C FULL EXPERIMENT RESULTS
We present the detailed experiment results on all LMM+LLM combinations in Table 4, from which the results of the best-performing combination is highlighted in Table 1. We restrict our experiments to two 7B LMMs and do not extend to the larger LLaVA1.6-13B, as its M-CoT results do not consistently outperform the 7B counterpart across most datasets, with improvements observed only on Harm-P and MultiOFF, while incurring substantially higher computational overhead. Notably, the 13B model shares the same vision encoder and differs primarily in language model capacity, offering limited visual advantage. We therefore focus on 7B LMMs as a more efficient choice aligned with our low-resource design objective. By default, each unimodal LLM reasons over visual information extracted solely by one paired 7B LMM. For FHM, we experiment with a specific collaborative setup between the two LMMs: Qwen2VL-7B is responsible for extracting disability-related visual cues, while LLaVA1.6-7B handles others. This design is motivated by Qwen2VL-7B’s lower tendency to hallucinate when identifying such information. This FHM-specific configuration is denoted as “LLaVA+Qwen” thereafter. In practice, practitioners can evaluate how well different LMMs’ (or their combinations’) visual understanding aligns with the target context based on sample memes to better determine which LMM setup offers better robustness in visual content recognition. Likewise, the selection of LLM+LMM combinations can be determined empirically according to pilot results.
As shown in Table 4, U-CoT+’s performance gains over U-CoT stay prominent across different LMM+LLM combinations in most datasets, indicating its overall effectiveness. Compared to more capable up-scaled LLMs, 7~8B LLMs struggle to achieve highly comparable results to that of state-of-the-art SFT baselines, and can only match or outperform GPT-4o-mini’s zero-shot results on FHM, HarMeme and MultiOFF. This suggests that, even when converted to a unimodal text-only inference setting, harmful meme detection remains reasoning-intensive, placing requirements on the general reasoning abilities of the underlying LLMs.
C.1. Meme2Text: From Multi- to Unimodal
C.1.1. Visual Information Extraction
We experiment with two widely used open-source 7B LMMs i.e., LLaVA1.6-Mistral-7B (Liu et al., 2023a) and Qwen2-VL-7B (Qwen2-VL, 2024). They have been pre-trained to gain strong capabilities in image understanding, embedded text recognition and multimodal inference, and are generally accessible under most computational resource conditions.
We identify various types of visual clues that are critical for downstream harmful meme detection and design the questions accordingly to prompt LMMs to extract the information through visual question answering (VQA). In particular, for the synthetic FHM dataset, since the visual style of its images is more like that of regular photographs other than that of authentic online memes, the captions of some memes do not directly relate to the visual contents. Therefore, we explicitly add the instruction “Ignore any overlaid text or caption.”, denoted by [Ignore], after the questions to force LMMs to focus solely on the visual content without being distracted by the overlaid captions. For most binary questions, we adopt this template for a unified output format: “Start your response with “Yes,” or “No,” before giving the explanation.”, denoted by [Fm]. The overlaid text or caption in a meme is denoted by [OCR]. Templates for visual questions:
  1. Human Is there any human subject in the given image? ([Ignore]) [Fm]
  2. #Human Does the image include more than one human subject? ([Ignore]) [Fm]
  3. Gender What is/are the gender(s) of the human subject(s) in the image?
  4. Race What is/are the perceived race(s) of the human subject(s) in the image?
  5. Appearances What are the distinctive physical appearance characteristics of the human subject(s) in the image?
  6. Disability Does any of the human subjects in the image have any disability?
  7. Celebrity (i) Who is/are the human subject(s) in the image? (ii) Is any celebrity or historical figure portrayed in the image? If yes, who are they? If no, just output “No.”.[Fm] (iii) FHM-specific: a. Does the image portray Adolf Hitler? If yes, output “Yes, Adolf Hitler is portrayed in the image.” Otherwise, if you are not sure about the identity of human subject, just output “No, I can’t tell.” b. Does the image portray Anne Frank, the Jewish girl who hid from the Nazis during World War II? If yes, output “Yes, Anne Frank is portrayed in the image.” Otherwise, if you are not sure about the identity of human subject, just output “No, I can’t tell.”
@@ 6. Disability Does any of the human subjects in the image have any disability?
  1. Adult Content Does the image’s visual content contain adult content?[Ignore][Fm]
  2. Female Is/Are there any human subject(s) in the image female?[Ignore][Fm]
  3. Sexual (i) Does the image’s visual content highlight the sexiness of the female’s figure in a way that is sexually provocative?[Ignore][Fm] (ii) Does this image highlight sexual body parts of the female subject(s), such as the breast, the hip/buttock, or the genital?[Ignore][Fm] (iii) Does/Do the female subject(s) in the image appear to be overweight?[Ignore][Fm] (iv) Does/Do the female subject(s) in the image appear to be of large body size (considered as fat)?[Ignore][Fm]